Based on this balanced equation:


2LiOH + H2S  → Li2S  + 2H2



How many moles of Li2S will be produced from 150.17 g of LiOH and excess H2S?


1
Expert's answer
2021-04-11T23:50:40-0400

Molar Mass of LiOH = 23.95

Molar Mass of H2S= 34.1

Molar Mass of Li2S = 45.95

From the equation,

= 2/3 × 150.17 = 100.1133g of LiOH

= 100.1133 /23.95 = 4.18moles of LiOH

1/3 × 150.17 = 50.0567g of H2S

50.0567/34.1 = 1.468moles of H2S

2 moles of LiOH = 1 mole of Li2S

= 4.18/2 = 2.09 moles

1 mole of H2S = 1 mole of Li2S

= 1 × 1.468 = 1.468 mole

= 1.468 + 2.09

= 3.558 moles of Li2S